# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ III-B Main theorem—Convergence rate of ProxSkip
ProxSkip $\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{\rho}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{\rho\epsilon}\right)$ $\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p\sqrt{\chi}}{\sqrt{\rho}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
@@
Achieving acceleration by $p$ and $n$. According to (7), (8), and (9), when $\epsilon$ is sufficiently small, the convergence rate is dominated by noise and is unaffected by the graph parameter $1-\lambda_{2}$ for ProxSkip. After enough transient time, ProxSkip with $\mathcal{O}\big{(}\frac{p\sigma^{2}}{n\epsilon^{2}}\big{)}$ (or $\tilde{\mathcal{O}}\big{(}\frac{p\sigma^{2}}{n\epsilon}\big{)}$ for the strongly convex setting) achieves linear speedup by the probability of communication $p$ and the number of nodes $n$.
@@
N-CVX/CVX: $\mathcal{O}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{3}}{{2}}}\right),$
S-CVX: $\tilde{\mathcal{O}}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{1}}{{2}}}\right).$
Here, $\varsigma^{2}$ represents the function heterogeneity constant such that $\nicefrac{{1}}{{n}}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x})-\nabla f(\mathbf{x}^{\star})\|^{2}\leq\varsigma^{2}$. We note that ProxSkip lacks the additional term $\frac{\varsigma}{\rho}\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}$ ($\frac{\varsigma}{\rho}\frac{1}{\sqrt{\epsilon}}$ for the strongly convex case). Thus, ProxSkip effectively eliminates dependence on the data heterogeneity level $\varsigma^{2}$.
Comparable with existing distributed algorithms incorporating local updates. When $p<\lambda_{2}$, we have $\chi=\mathcal{O}(\max\{1,\frac{1-p}{1-\lambda_{2}}\})=\mathcal{O}(\frac{1}{1-\lambda_{2}})$; when $p>\lambda_{2}$, $\chi=\mathcal{O}(\max\{1,\frac{1-p}{1-\lambda_{2}}\})=\mathcal{O}(1)$. Letting $p=\nicefrac{{1}}{{K}}$ and highlighting the network quantities, the second + third term of the communication complexity of ProxSkip $\mathcal{O}\left(K^{-1}\rho^{-1}+\rho^{-2}\right)$ when $p<1-\rho$, $\mathcal{O}\left(K^{-1}\rho^{-\nicefrac{{1}}{{2}}}+\rho^{-1}\right)$ when $p\geq 1-\rho$. Compared with GT based methods [35, 36, 37], the network dependent bounds are improved. Since the first item of communication complexity for these algorithms is the same, the convergence rates of ProxSkip are comparable with these existing distributed algorithms incorporating local updates.
@@ III-C Linear speedup with network-independent stepsizes
Suppose that Assumptions 1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. If $0<\alpha\leq\frac{1}{2L}$, $\beta=p$, there exists $\chi=\mathcal{O}(\max\{\frac{1}{p},\frac{1}{1-\lambda_{2}},\frac{1-p}{1-\lambda_{2}}\})$ such that
@@
Notably, [51] for the first time prove that NIDS/ED/D2 can achieve linear speedup with network-independent stepsizes. However, it remains an open question whether, with local updates, a linear speedup bound can be achieved using network-independent stepsizes. Theorem 2 offers a positive response to this question.
III-D Proof sketch of the main theorem
The existing convergence analysis of ProxSkip [43, 44, 45, 47, 48, 49] relies on primal-dual methodologies. Nevertheless, these analyses are limited to the use of first-order (stochastic) gradient information, leading to a suboptimal exploitation of the available function data. We propose a new proof that, in order to fully utilize function and gradient information, we use matrix factorization techniques to equivalently transform the iteration of ProxSkip into “SGD + consensus” form. Here, we provide a proof sketch for Theorem 1 concerning non-convex objectives. The complete proof of Theorem 1 and Corollary 1 can be found in the Appendix. Here, we provide a proof sketch for Theorem 1 concerning non-convex objectives.
Step 1. (Lemma 1) We first give the equivalent form of update (2) as follows.
$\underbrace{\bar{{\bf{x}}}^{t+1}=\bar{{\bf{x}}}^{t}-\alpha\bar{{\bf{g}}}^{t}}_{\mathrm{SGD}},\quad\underbrace{\mathcal{E}^{t+1}={\bf{\Gamma}}\mathcal{E}^{t}+\bm{\Theta}_{1}^{t}+\bm{\Theta}_{2}^{t}}_{\mathrm{consensus}},$ (11)
where $\bar{{\bf{g}}}^{t}=\frac{1}{n}\sum_{i=1}^{n}{\bf{g}}_{i}^{t}$, $\|{\bf{\Gamma}}\|<1$, $\mathcal{E}^{t}$ measures “consensus”, $\bm{\Theta}_{1}^{t}$ is related to the stochastic gradient, and $\bm{\Theta}_{2}^{t}$ measures “communication error” ($\bm{\Theta}_{2}^{t}=0$, if $\theta_{t}=1$). This description may be less rigorous, but it helps to understand the proof more clearly. See Lemma 1 in Appendix for more details.
Step 2. (Lemma 2) Based on this equivalent update of ProxSkip and by the $L$-smoothness of $f_{i}$, we establish the following descent inequality.
$\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t+1})\right]\leq$ $f(\bar{{\bf{x}}}^{t})-\frac{\alpha}{2}\big{\|}\nabla f(\bar{{\bf{x}}}^{t})\big{\|}^{2}$
$+\frac{2\alpha L^{2}}{n}\|\mathcal{E}^{t}\|^{2}+{\frac{L\alpha^{2}\sigma^{2}}{2n}}.$
Taking average for both sides over $t=0,1,\ldots,T-1$, we have that
$\leq\frac{2F_{0}}{\alpha T}+\frac{4L^{2}}{nT}\sum_{t=0}^{T-1}\|\mathcal{E}^{t}\|^{2}+{\frac{\alpha L\sigma^{2}}{n}}.$ (12)
Step 3. (Lemma 2) We bound $\mathbb{E}\!\left[\|{\bf{\Gamma}}\mathcal{E}^{t}+\bm{\Theta}_{1}^{t}\|^{2}\right]$ and $\mathbb{E}\!\left[\|\bm{\Theta}_{2}^{t}\|^{2}\right]$ by $\|\mathcal{E}^{t}\|^{2}$, $\sigma^{2}$, and $\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$.
Step 4. (Lemma 2) Subsequently, we establish the following consensus inequality.
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|^{2}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|^{2}+A_{1}n\alpha^{4}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}$
$+A_{2}n\alpha^{2}\sigma^{2}+A_{3}\alpha^{4}\sigma^{2},$
where $10<\tilde{\gamma}<$ 1 and $A_{1},A_{2},A_{3}>0$. Unrolling this recurrence, we have
$\mathbb{E}\!\left[\|\mathcal{E}^{t}\|^{2}\right]\leq$ $\frac{\|\mathcal{E}^{0}\|^{2}}{1-\tilde{\gamma}}+\frac{A_{2}n\alpha^{2}\sigma^{2}+A_{3}\alpha^{4}\sigma^{2}}{1-\tilde{\gamma}}$
$+A_{1}n\alpha^{4}\sum_{k=0}^{t-1}\tilde{\gamma}^{t-1-k}\|\nabla f(\bar{{\bf{x}}}^{k})\|^{2}.$
Step 5. Since
$\sum_{t=0}^{T-1}\sum_{k=0}^{t-1}\tilde{\gamma}^{t-1-k}\|\nabla f(\bar{{\bf{x}}}^{k})\|^{2}\leq\frac{1}{1-\tilde{\gamma}}\sum_{t=0}^{T-1}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2},$
letting $\alpha$ such that
$\frac{4L^{2}A_{1}\alpha^{4}}{1-\tilde{\gamma}}\leq\frac{1}{2},$
it gives that
$\frac{4L^{2}}{nT}\sum_{t=0}^{T-1}\big{\|}\mathcal{E}^{t}\big{\|}_{\mathrm{F}}^{2}\leq\mathcal{O}(\alpha^{2}\sigma^{2})+\frac{1}{2T}\sum_{t=0}^{T-1}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}.$
Combining it with (III-D), we complete the proof, i.e.,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right).$
@@ V Conclusion
I Preliminaries
I-A Basic Facts
@@
$\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq D_{f}(x,y),\text{ and }\langle\nabla f(x)-\nabla f(y),x-y\rangle\geq D_{f}(y,x).$ (13)
@@ I-B Notations
communication noise: ${\bf{E}}^{t}=\frac{\theta_{t}-1}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}.$
@@
Therefore, the matrix ${\bf{W}}_{a}$ and ${\bf{W}}_{b}$ can be decomposed as
${\bf{W}}_{a}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}{\bf{1}}&0\\ 0&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{a}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],~{}{\bf{W}}_{b}^{2}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}0&0\\ 0&\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{b}^{2}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],$ (27)
where $\hat{{\bf{\Lambda}}}_{a}={\bf{I}}-\frac{1}{2\chi}({\bf{I}}-\hat{{\bf{\Lambda}}})$, $\hat{{\bf{\Lambda}}}_{b}=\sqrt{{\bf{I}}-\hat{{\bf{\Lambda}}}}$. Since $\lambda_{i}\in(-1,1)$ for $i=2,\dots,n$, it holds that $1-\frac{1}{2\chi}(1-\lambda_{i})\in[0,1)$ and $0\preceq{\bf{W}}_{a}\prec{\bf{I}}$ for $\chi\geq 1$.
II Proof of Theorem 1 and Corollary 1
II-A Transformation and Some Descent Inequalities
@@
$\mathcal{E}^{t+1}$ $=\underbrace{{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{1}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]}_{:=\mathbb{G}^{t}}+\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right]}_{:=\mathbb{F}^{t}},$ (28f)
where $\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1$,
@@
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{(1+\lambda_{n})(1-\lambda_{2})}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+(1-p))}{\chi^{2}},$ (31)
@@
$\tilde{\gamma}=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{1-\lambda_{2}}}{1-\gamma}+\frac{2(1-p)\big{(}3+\frac{24\chi\alpha^{2}L^{2}}{(1+\lambda_{n})(1-\lambda_{2})}\big{)}}{\chi^{2}}.$ (32)
@@
Proof See Appendix V. $\Box$
II-B Convergence Analysis: Non-convex

### v4

@@ III-B Main Theorem—Convergence Rate of ProxSkip
ProxSkip $\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p\chi}}{\sqrt{\rho}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{\rho\epsilon}\right)$ $\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{\sqrt{p\chi}}{\sqrt{\rho}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
@@
Achieving acceleration by $p$ and $n$. According to (10), (11), and (12), when $\epsilon$ is sufficiently small, the convergence rate is dominated by noise and is unaffected by the graph parameter $1-\lambda_{2}$ for ProxSkip. After enough transient time, ProxSkip with $\mathcal{O}\big(\frac{p\sigma^{2}}{n\epsilon^{2}}\big)$ ($\tilde{\mathcal{O}}\big(\frac{p\sigma^{2}}{n\epsilon}\big)$ for SC) achieves linear speedup by the probability of communication $p$ and the number of nodes $n$. Note that reducing the communication probability $p$ beyond a certain point is counterproductive. Continuously reducing the communication probability $p$ cannot indefinitely lower the communication complexity, as at least $\mathcal{O}(\frac{\chi}{\rho\epsilon})$ ($\frac{\chi}{\rho}\mathrm{log}\nicefrac{{1}}{{\epsilon}}$ for SC) communication rounds are required. Moreover, since the stepsize $\alpha\leq\mathcal{O}(\frac{p(1-\lambda_{2})}{\chi L})$ ($\alpha\leq\mathcal{O}(\frac{p\mu(1-\lambda_{2})}{\chi L^{2}})$ for SC), an excessively small $p$ not only fails to improve communication efficiency but also forces a tiny stepsize, which severely hampers the convergence rate.
@@
N-C/C: $\mathcal{O}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{3}}{{2}}}\right),$
SC: $\tilde{\mathcal{O}}\left(\left(\nicefrac{{\sigma}}{{\sqrt{\rho K}}}+\nicefrac{{\varsigma}}{{\rho}}\right)\epsilon^{-\nicefrac{{1}}{{2}}}\right).$
Here, $\varsigma^{2}$ represents the function heterogeneity constant such that $\nicefrac{{1}}{{n}}\sum_{i=1}^{n}\|\nabla f_{i}(\mathbf{x})-\nabla f(\mathbf{x}^{\star})\|^{2}\leq\varsigma^{2}$. We note that ProxSkip lacks the additional term $\frac{\varsigma}{\rho}\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}$ ($\frac{\varsigma}{\rho}\frac{1}{\sqrt{\epsilon}}$ for SC). Thus, ProxSkip effectively eliminates dependence on the data heterogeneity level $\varsigma^{2}$.
Comparable with existing distributed algorithms incorporating local updates. When $p<\lambda_{2}$, we have $\chi=\mathcal{O}(\frac{1}{1-\lambda_{2}})$, the communication complexity of ProxSkip is
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{1-\lambda_{2}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{(1-\lambda_{2})^{2}}\frac{1}{\epsilon}\right).$
When $p>\lambda_{2}$, we have $\chi=\mathcal{O}(1)$, the communication complexity is
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{p}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{1-\lambda_{2}}\frac{1}{\epsilon}\right).$
Letting $p=\nicefrac{{1}}{{K}}$ and highlighting the network quantities, the second + third term of the communication complexity of ProxSkip $\mathcal{O}\left(K^{-\nicefrac{{1}}{{2}}}\rho^{-1}+\rho^{-2}\right)$ when $p<1-\rho$, $\mathcal{O}\left(K^{-\nicefrac{{1}}{{2}}}\rho^{-\nicefrac{{1}}{{2}}}+\rho^{-1}\right)$ when $p\geq 1-\rho$. Compared with GT based methods [35, 36, 37], the network dependent bounds are improved. Since the first item of communication complexity for these algorithms is the same, the convergence rates of ProxSkip are comparable with these existing distributed algorithms incorporating local updates.
@@ III-C Linear Speedup With Network-Independent Stepsizes
Suppose that Assumptions 1, 2, and 4 hold, and $f_{i}$ is $\mu$-strongly convex for some $0<\mu\leq L$. If $0<\alpha\leq\frac{1}{2L}$, $\beta=p$, there exists $\chi=\mathcal{O}(\max\{1,\frac{1}{1-\lambda_{2}},\frac{1-p}{1-\lambda_{2}}\})$ such that
@@ V Conclusion
VI Preliminaries
VI-A Basic Facts
@@
[…]
@@ VI-B Notations
communication noise: ${\bf{E}}^{t}=\frac{\theta_{t}-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}.$
@@
$\hat{{\bf{P}}}^{\sf T}\hat{{\bf{P}}}={\bf{I}},\ {\bf{1}}^{\sf T}\hat{{\bf{P}}}=0,\ \hat{{\bf{P}}}\hat{{\bf{P}}}^{\sf T}={\bf{I}}-\frac{1}{n}{\bf{11}}^{\sf T}.$
Let $\widehat{{\bf{W}}}_{a}:=\mathbf{I}-\frac{p}{2\chi}{\bf{W}}_{b}^{2}$. Therefore, the matrix $\widehat{{\bf{W}}}_{a}$ and ${\bf{W}}_{b}$ can be decomposed as
$\widehat{{\bf{W}}}_{a}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}{\bf{1}}&0\\ 0&\hat{{\bf{\Lambda}}}_{a}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{a}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],~{\bf{W}}_{b}^{2}=\left[\begin{array}[]{cc}{\bf{1}}&\hat{{\bf{P}}}\\ \end{array}\right]\underbrace{\left[\begin{array}[]{cc}0&0\\ 0&\hat{{\bf{\Lambda}}}_{b}^{2}\\ \end{array}\right]}_{:={\bf{\Lambda}}_{b}^{2}}\left[\begin{array}[]{c}\frac{1}{n}{\bf{1}}^{\sf T}\\ \hat{{\bf{P}}}^{\sf T}\\ \end{array}\right],$ (28)
@@
VII Proof of Theorem 1 and Corollary 2
VII-A Transformation and Some Descent Inequalities
@@
$\mathcal{E}^{t+1}$ $=\underbrace{{\bf{\Gamma}}\mathcal{E}^{t}-\alpha{\bf{Q}}^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}^{2}\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{1}^{t}+\hat{{\bf{P}}}^{\sf T}{\bf{\Sigma}}_{2}^{t}\\ \end{array}\right]}_{:=\mathbb{G}^{t}}+\underbrace{{\bf{Q}}^{-1}\left[\begin{array}[]{c}-\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{E}}^{t}\\ \end{array}\right]}_{:=\mathbb{F}^{t}},$ (29f)
where $\gamma\triangleq\|{\bf{\Gamma}}\|=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}<1$, ${\bf{E}}^{t}=\frac{\theta_{t}-p}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$,
@@
$\|{\bf{Q}}\|^{2}\leq 2\text{ and }\|{\bf{Q}}^{-1}\|^{2}\leq\frac{2\chi}{p(1+\lambda_{n})(1-\lambda_{2})}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]\leq$ $\tilde{\gamma}\|\mathcal{E}^{t}\|_{\mathrm{F}}^{2}+\frac{4n\alpha^{4}L^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}+\frac{2\alpha^{4}L^{2}\sigma^{2}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2n\alpha^{2}\sigma^{2}(2\chi^{2}+p(1-p))}{\chi^{2}},$ (32)
@@
$\tilde{\gamma}=\gamma+\frac{32\alpha^{2}L^{2}+16\alpha^{4}L^{4}\frac{2\chi}{p(1-\lambda_{2})}}{1-\gamma}+\frac{2p(1-p)\big(3+\frac{24\chi\alpha^{2}L^{2}}{p(1+\lambda_{n})(1-\lambda_{2})}\big)}{\chi^{2}}.$ (33)
@@
Proof See Appendix X. $\Box$
VII-B Convergence Analysis: Non-convex